OpenStudy (anonymous):

I tried to solve this as many different ways as i could come up with but none of them were any of the provided answers below..can someone please show me how to answer this? Evaluate the expression for the given value of the variable(s). -x^2 - 4x; x = –3 A. –17 B. 11 C. 3 D. –1

OpenStudy (anonymous):

\[y(x)= -x^2-4x \ \quad ; \quad x=-3\] \[y(-3) = -(-3)^2-4(-3)\] \[=-9 +12\] \[ = 3\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

^^ welcome